Ifx= root 3+1 upon 2, find the value of 4xcube + 2xsquare-8x+7. pls solve question RD sharma chapter-3 question-11

Grade:9

1 Answers

pankaj kumar raman
11 years ago

x=√3+1/2

x-1/2=√3

squaring on both sides

(x-1/2)²=3

4x²+1-x=3

4x²-x-2=0......................(1)

take the lhs and divide it by 4xcube +2x² -8x+7

hence

4xcube +2x² -8x + 7= (4x²-x-2)(x+3/4)+(-21/4x+17/2)

from eq (1)

                             =0(x+3/4) + (-21/4x+17/2)

                             =-21/4(√3+1/2) + 17/2

                             =47-42√3 / 8 ans.